OpenStudy (anonymous):

How much heat is contained in 100 kg of water at 60°C?

OpenStudy (shamim):

Can u tell me the heat needed for 100 kg water frm 0k to 60 degree celcius

OpenStudy (shamim):

60 degree celcius=60+273=333k

OpenStudy (shamim):

0k=-273 degree celcius Right?

OpenStudy (shamim):

Heat needed for 100 kg water frm -273 degree celcius to 0 degree celcius Q1=ms(theta2-theta1) Q1=100*2100*(0--273)=100*2100*273=?J

OpenStudy (shamim):

* not water but ice

OpenStudy (shamim):

Now heat needed for 100 kg ice frm 0degree celcius to water of 0 degree celcius Q2=mLf=100*Lf=?J

OpenStudy (shamim):

Heat needed for 100 kg water frm 0 degree celcius to 60 degree celcius Q3=ms(theta2-theta1) Q3=100*4200*(60-0)=?

OpenStudy (shamim):

Q=Q1+Q2+Q3=?

OpenStudy (shamim):

Response plz!!!!!!

OpenStudy (shamim):

M i right????